OpenStudy (anonymous):

Is the series Σ(1/log(n)) , for n=2 -> infinity convergent and if yes how to prove it?

OpenStudy (anonymous):

definitely not convergent

OpenStudy (anonymous):

for one thing \(\ln(n)<n\) making \(\frac{1}{\ln(n)}>\frac{1}{n}\) so comparison test will give you a quick answer, but by eyeballing it you can tell it does not converge because log grows very slowly

OpenStudy (anonymous):

Thanks!